Solve the equations graphically if possible:
{ x + 3y = -6
{ 2x - 4y = -12



{ x = 4 + y
{ y = 3 + x

Answer by ewatrrr(24785)

 
Hi
Re: determining the coordinates to plot:
1)Letting x = 0 and y = 0 to determine those ordered pairs is very common
(as done in the following: for ex. finding (0,-2)and (-6,0)
OR
2)One may 'plug-in' any value of x to determine corresponding y-value
Most'plug-in' an 'easy' x-value..may even be a fraction...if that results in an 'easy' y
x + 3y = -6 Points (0,-2) and (-6,0) on this line (Green)
2x - 4y = -12 Points (0,3) and (-6,0) on this line Blue
Plot points, connect with their respective lines.
Point of intersection (-6,0): this ordered pair is the solution for this system

x = 4 + y OR y = x-4
y = 3 + x OR y = x+3